Calculate the activation energy for a second-order reaction if its rate constant in- creases from 2.5 mol/(L·s) to 250 mol/(L·s) as the temperature increases from 298 to 340 K. 


1
Expert's answer
2021-10-27T06:41:18-0400

We are given that:

When T1

​=298K

Let k1

​=k

When T

=340K

k2

​=2k

Substituting these values the equation:

"Log\frac{2K}{K}" = "\frac{E_a}{2.303×8.314}" ("(\frac{340-298}{340×398})"

Ea135320=804.18

Ea=5.94×10-3J